OpenStudy (anonymous):

Which equation below is not an equation that represents the graph shown below? y + 2 = -3(x - 1) 3x + y =1 y - 1 = -3(x + 2) y = -3x + 1

OpenStudy (anonymous):

 The line looks like to be right on the 1 one the y-axis.

OpenStudy (anonymous):

\(y - 1 = -3(x + 2) \) is the wrong one

OpenStudy (anonymous):

if we multiply out we get \[y-1=-3x-6\] and then \[y=-3x-5\] which is incorrect. all the others are the same, they are all \[y=-3x+1\]